Please help me solve this problem:
How much interest does a $10,000 investment earn at 6% over 15 years compounded quarterly?

Answer by ewatrrr(24785)


Hi
Note: A= P(1+r/n) ^nt OR I + P = P(1+r/n) AND I =P[(1+ r/n)^nt - 1]
I = $10,000[(1+ .06/4)^4*15 - 1] = $10,000[(1.015^60 - 1] = $14,432.20
CHECKING our Answer***
A= P(1+r/n) ^nt
$24,432.20 = 10,000(1.015)^60